OpenStudy (anonymous):

hey how do i get the square root of my answer in a pythagoreaon identity question. for example for my answer i got cos^2t=24/49

OpenStudy (anonymous):

so cos t = sqrt (24/49)

OpenStudy (anonymous):

im looking to know if it can be square rooted the question asks for a square and m wondering if how can i do that with a fraction

OpenStudy (anonymous):

ofc it can be

OpenStudy (anonymous):

wat lol

OpenStudy (anonymous):

sqrt(24/49) = sqrt(24)/sqrt(49)

OpenStudy (anonymous):

so i have to square each number individually

OpenStudy (anonymous):

wat would be the answer

OpenStudy (anonymous):

cause on my example they are getting decimals

OpenStudy (anonymous):

cos t = +or- 0.6998

OpenStudy (anonymous):

and cos t cannot be negative so answer is approx 0.7

OpenStudy (anonymous):

how did u get that

OpenStudy (anonymous):

using calculator

OpenStudy (anonymous):

ok thanks lol dont have one of those fancy calculators lol

OpenStudy (anonymous):

wats ur name r u gonna be on here for awhile

OpenStudy (anonymous):

you have any calculator that can calculate the sqrt of a number ?

OpenStudy (anonymous):

not at the moment

OpenStudy (anonymous):

use the windows calculator then :)

OpenStudy (anonymous):

i found one thanks